All ISS sites operational. The new He adapter was received and the regulators switched out. Balloon filling now takes less than 3 min. vs. 15-16 min. with other regulator, much to the relief of sounding operators. Started with 30 cu. ft. and increasing until desired ascent rates reached. Looks like value will be between 40-45 cu. ft. Moved the temperature logger from the sounding trailer to the wind profiler shed to monitor/log temperatures there. Gary has some concern about the high values seen in the recorded CPU temperatures and would like to correlate them with temperatures inside the shed. Weather has cooled some so not expecting any thermal issues in the next few days. Mentioned at daily meeting that the 4 lidars at the upper orange site are all on the same circuit breaker and to consider moving loads to multiple breakers when convenient. No input received.
Isabel and Kate up to speed on soundings with Gary learning.